# Interpretable Failure Analysis in Multi-Agent Reinforcement Learning Systems

### v1

@@ Abstract.
Multi-Agent Reinforcement Learning (MARL) is increasingly deployed in safety-critical domains, yet methods for interpretable failure detection and attribution remain underdeveloped. We introduce a two-stage gradient-based framework that provides interpretable diagnostics for three critical failure analysis tasks: (1) detecting the true initial failure source (Patient-0); (2) validating why non-attacked agents may be flagged first due to domino effects; and (3) tracing how failures propagate through learned coordination pathways. Stage 1 performs interpretable per-agent failure detection via Taylor-remainder analysis of policy-gradient costs, declaring an initial Patient-0 candidate at the first threshold crossing. Stage 2 provides validation through geometric analysis of critic derivatives—first-order sensitivity and directional second-order curvature aggregated over causal windows to construct interpretable contagion graphs. This approach explains “downstream-first” detection anomalies by revealing pathways that amplify upstream deviations. Evaluated across 500 episodes in Simple Spread (3 and 5 agents) and 100 episodes in StarCraft II using MADDPG and HATRPO, our method achieves 88.2–99.4% Patient-0 detection accuracy while providing interpretable geometric evidence for detection decisions. By moving beyond black-box detection to interpretable gradient-level forensics, this framework offers practical tools for diagnosing cascading failures in safety-critical MARL systems.
@@ 1. Introduction
Multi-Agent Reinforcement Learning (MARL) has emerged as a pivotal framework for coordinating autonomous systems across safety-critical domains including robotic fleets, transportation systems, communication networks, and power grids Petrenko and Gurchinskiy (2021); Bakakeu et al. (2020); Xu et al. (2022); Panfili et al. (2018); Peng and Shen (2021); Noureddine et al. (2017). In these settings, where multiple learned policies interact, a single local deviation can ripple through the entire team causing other agents to fail Liu et al. (2022).
@@
Despite these needs, explainable MARL (xMARL) remains underexplored. Shapley-value attribution Heuillet et al. (2022); Li et al. (2021) has been used to quantify each agent’s contribution to team reward in cooperative scenarios. Although such attribution offers insights into credit assignment, it does not reveal inter-agent dynamics or cascading failures. Complementary work Boggess et al. (2023) provides policy-level contrastive explanations by checking user-specified temporal task queries against abstracted policies via probabilistic model checking. While valuable, these approaches are not designed to analyze failure dynamics in team settings, such as how failures unfold, which agent was the inciting point of failure, and why.
@@
In contrast, single-agent Explainable RL (XRL) has recently shifted toward safety debugging and testing, using explanations to surface vulnerabilities and guide falsification (e.g., Shefin et al. (2025); Le et al. (2025)). For instance, xSRL Shefin et al. (2025) combines local post-hoc explanations with global abstract policy graphs to identify and patch safety violations. AEGIS-RL Le et al. (2025) builds on xSRL by developing a verification approach that integrates formal model checking with risk- and uncertainty-guided falsification. However, repurposing these single-agent methods for MARL is not straightforward due to the heightened complexity of multi-agent systems. The explainability challenge in MARL surpasses that of single-agent RL, primarily due to the inherent coupling and dependencies among agents. While single-agent explanations can focus on an individual policy, MARL requires untangling a web of interdependencies, compounded by emergent collective behaviors and environmental non-stationarity. These dynamic feedback loops obscure cause-effect relationships, making it difficult to distinguish between agent-level and system-level failures.
@@
Moreover, conventional failure analysis in MARL often tracks global metrics such as team reward or derived statistics over time Abdallah and Lesser (2008); Xu et al. (2023). Such global performance metrics can obscure policy instability in individual agents Abdallah (2009); Khlifi et al. (2023) as we show in our experiments. Moreover, they neither identify the first agent to enter a non-robust state nor reveal how deviations propagate, all of which are needed to pinpoint weak points and pathways for proactive mitigation.
@@
Therefore, this paper formalizes the interpretable failure analysis problem in MARL around three core questions; Q1. Who is the true Patient-0: the first agent to enter a non-robust state?, Q2. Why might a non-attacked agent be flagged first? How do we distinguish true sources from amplified downstream effects?, and Q3. How does instability propagate across agents over time through the system’s learned coordination pathways? We propose to address these questions through a two-stage, gradient-based forensic method. Stage 1 performs per-agent local detection by probing the Taylor-error of a policy-gradient cost; the earliest threshold crossing declares an initial Patient-0 candidate (Q1). Stage 2 validates or corrects this candidate by tracing accelerating upstream influence using critic derivatives—first- and second-order action sensitivities aggregated over causal windows (Q2). This approach naturally yields constructions of directed contagion graphs that summarize propagation pathways, influence strength, and timing (Q3).
@@ 2. Related Work
Explainability in MARL (xMARL). Research in explainable MARL is limited and primarily splits into (i) credit assignment Heuillet et al. (2022); Li et al. (2021) and (ii) specification- or query-driven explanations Boggess et al. (2023). Shapley-based methods Heuillet et al. (2022); Li et al. (2021) quantify each agent’s contribution to the team reward, offering insights into fairness and cooperation but not into failure dynamics or inter-agent cascades . Recent work continues to advance credit assignment, such as Multi-level Advantage Credit Assignment (MACA) Zhao and Xie (2025), which performs counterfactual reasoning across different levels of agent coordination but remains focused on optimizing training rather than diagnosing post-hoc failures . Complementarily, policy-level contrastive explanations Boggess et al. (2023) verify temporal task queries (e.g., encoded in PCTL*) via probabilistic model checking on an abstracted policy . While both lines illuminate aspects of MARL behavior, neither targets post-hoc failure analysis under compromised agents: they do not identify the entry point of a failure nor reconstruct how influence propagates across agents over time. Our method addresses this gap by detecting the first non-robust agent and mapping directed, time-resolved propagation paths using centralized-critic sensitivities.
Single-agent XRL and Policy Summaries. A rich body of XRL provides explanations for a single agent, including policy summarization via abstract policy graphs Amir and Amir (2018); McCalmon et al. (2022), contrastive explaination methods van der Waa et al. (2018), and frameworks that integrate local and global summaries Shefin et al. (2025); Le et al. (2025). Safety-focused XRL (e.g., xSRL Shefin et al. (2025)) ties explanations to debugging and patching unsafe behavior but remains single-agent in scope. Directly porting these methods to MARL misses the core challenge: the coupling and non-stationarity among agents obscure cause-effect relations across the team. Our work departs by explicitly modeling cross-agent influence during failure episodes and delivering a cascade map that explains how a local deviation spreads at run time.
Attacks and Robustness. The vulnerability of cooperative MARL to adversarial pressure is well-studied, including attacks that degrade team performance Liu et al. (2022); Kazari et al. (2023a); Li et al. (2023); Lischke et al. (2022). These works establish vulnerabilities and propose defenses, but generally focus on performance degradation or anomalous behavior detection without explaining the failure’s origin and propagation pathway. Our contribution is orthogonal and complementary: a post-hoc explainer that identifies Patient-0 and attributes influence across agents.
@@ 3. Approach
We propose a two-stage, interpretable gradient-based method for MARL failure forensics. The key insight is that a failure typically begins with one agent and then propagates via learned coordination. Stage 1 performs per-agent local detection by probing policy-gradient Taylor error and declares Patient-0 as the first threshold crossing. Stage 2 validates or corrects it by tracing accelerating upstream influence using critic derivatives—first-order sensitivity and directional curvature—aggregated over a short causal window. These scores yield a directed contagion graph that summarizes who influenced whom, the strength and frequency of amplification, and how quickly deviations spread. We assume a single failure entry per episode, with subsequent failures arising from cascades through cooperative coupling. [^1] Relaxing the single–entry assumption is left to future work.
@@ 3.1. Two-Stage Patient-0 Detection and Validation
so that for small $\eta_{i}$, $\mathcal{L}_{i}^{t}(\eta_{i})\approx\tfrac{1}{2}\eta_{i}^{\top}\nabla_{o_{i}}^{2}J_{i}(o_{i}^{t},\tau_{i})\eta_{i}$. This provides a curvature probe in observation space without forming the full Hessian, similar to the Second-Order Identification of Non-Robust Directions test from single-agent RL Korkmaz and Brown-Cohen (2023). Large deviations indicate the policy is in a non-robust, high-curvature region where small observation changes could cause large policy shifts.
@@
In MARL algorithms with action-value critic (e.g., MADDPG Lowe et al. (2017)), we leverage the trained per-agent critic $Q_{i}(s,a_{1},\dots,a_{n})$ that conditions on the global state (or joint observations) and the joint action; this object encodes inter-agent dependencies learned during training. In algorithms that only learn a state-value function (e.g., HATRPO Kuba et al. (2022)), we fit a light post-hoc probe critic $\tilde{Q}_{i}(s,a_{1},\dots,a_{n})$ on frozen rollouts for attribution only (not for control), by regressing to TD or Monte-Carlo targets. [^2] Either critic is required only to be differentiable with respect to each action argument so that we can take first- and second-order derivatives. For discrete actions we use a standard continuous relaxation (e.g., softmax with temperature or Gumbel–Softmax during probing); for continuous actions we directly backpropagate. This yields well-defined partials $\partial Q_{i}/\partial a_{j}$ and $\partial^{2}Q_{i}/\partial a_{j}^{2}$ that quantify how agent $j$’s action locally perturbs agent $i$’s value and, by extension, its policy’s stability signals. We define the critic cost function as $L_{Q_{i}}=-Q_{i}$, thereby treating higher values as worse and lower values as better, which preserves the standard characteristics of a cost function.
@@ 4.1. Setup
Environments and algorithms. We use Simple Spread (cooperative navigation) with $n\!\in\!\{3,5\}$ agents and SMAC (Starcraft II). In Simple Spread, $n$ agents must navigate to $L$ landmarks through coordinated movement. Agents observe relative positions of other agents and landmarks, and receive shared rewards to minimize distance to landmarks. SMAC is a decentralized micromanagement benchmark where 3 Marine agents must defeat 3 Zealot opponents through coordinated focus fire and positioning. Each agent receives partial observations and controls one Marine unit, requiring emergent coordination strategies to overcome the enemy’s statistical advantages. We tested two MARL algorithms; MADDPG Lowe et al. (2017) and HATRPO Kuba et al. (2022).
@@
Failure Simulation/ Intervention Protocol. We simulate failures using the worst-action attack from adversarial MARL Li et al. (2023); Liu et al. (2022). We then created an intervention protocol to test our proposed work. For each ordered pair $(j\!\to\!i)$, we intervene at two classes of timesteps computed from the trained critic(s): Critical states where $G_{ij}$ is high and $D_{ij}\!>\!0$, and Robust states where $G_{ij}$ is low and $D_{ij}\!\leq\!0$. We replace $a_{j}$ with the same-strength attack action in both conditions and compare outcomes over an identical window $\mathcal{W}$. Per environment this yields $500$ base episodes $\times$ ordered pairs $\times 2$ conditions: 6,000 variants for n=3n{=}3 (6 pairs), and 20,000 for n=5n{=}5 (20 pairs). For the SMAC 3s_vs_3z, we experimented with $100$ base episodes, yielding 1,200 variants in total.
@@
Evaluation Metrics. To validate (Patient-0 identification), we report four accuracy measures: Stage-1 accuracy $\Pr(\hat{p}_{\mathrm{S1}}{=}p^{\star})$, Traceback accuracy $\%(\text{chain ends at }p^{\star})$, Combined accuracy after Stage-2 correction, and Correction rate (the fraction of downstream-first misidentifications successfully corrected by Stage-2, calculated over all downstream-first cases). We didn’t report Patient-0 baselines because prior works address different problems: Lischke et al. (2022) detects per-timestep anomalies using LSTMs on state-action traces without identifying the initial failure source, while Kazari et al. (2023b) develops decentralized GRU-based normality scoring for anomaly detection but does not identify the first inciting agent.
@@ 4.2. Results
Case 2 — Downstream-first flag with traceback correction (Simple Spread, n=5n{=}5, MADDPG). An adversarial perturbation is applied to agent 3 over $t{=}5\text{--}8$. Yet Stage 1 raises the first alarm at agent 1 at t=7t{=}7, creating a downstream-first cascade and a false Patient-0.

### v2

@@ Abstract.
Multi-Agent Reinforcement Learning (MARL) is increasingly deployed in safety-critical domains, yet methods for interpretable failure detection and attribution remain underdeveloped. We introduce a two-stage gradient-based framework that provides interpretable diagnostics for three critical failure analysis tasks: (1) detecting the true initial failure source (Patient-0); (2) validating why non-attacked agents may be flagged first due to domino effects; and (3) tracing how failures propagate through learned coordination pathways. Stage 1 performs interpretable per-agent failure detection via Taylor-remainder analysis of policy-gradient costs, declaring an initial Patient-0 candidate at the first threshold crossing. Stage 2 provides validation through geometric analysis of critic derivatives-first-order sensitivity and directional second-order curvature aggregated over causal windows to construct interpretable contagion graphs. This approach explains “downstream-first” detection anomalies by revealing pathways that amplify upstream deviations. Evaluated across 500 episodes in Simple Spread (3 and 5 agents) and 100 episodes in StarCraft II using MADDPG and HATRPO, our method achieves 88.2–99.4% Patient-0 detection accuracy while providing interpretable geometric evidence for detection decisions. By moving beyond black-box detection to interpretable gradient-level forensics, this framework offers practical tools for diagnosing cascading failures in safety-critical MARL systems.
@@ 1. Introduction
Multi-Agent Reinforcement Learning (MARL) has emerged as a pivotal framework for coordinating autonomous systems across safety-critical domains including robotic fleets, transportation systems, communication networks, and power grids petrenko2021multi; cps; Minrui; Panfili; Peng; Noureddine2017MultiagentDR. In these settings, where multiple learned policies interact, a single local deviation can ripple through the entire team causing other agents to fail MARLPrima.
@@
Despite these needs, explainable MARL (xMARL) remains underexplored. Shapley-value attribution shap; shap2 has been used to quantify each agent’s contribution to team reward in cooperative scenarios. Although such attribution offers insights into credit assignment, it does not reveal inter-agent dynamics or cascading failures. Complementary work temporal provides policy-level contrastive explanations by checking user-specified temporal task queries against abstracted policies via probabilistic model checking. While valuable, these approaches are not designed to analyze failure dynamics in team settings, such as how failures unfold, which agent was the inciting point of failure, and why.
@@
In contrast, single-agent Explainable RL (XRL) has recently shifted toward safety debugging and testing, using explanations to surface vulnerabilities and guide falsification (e.g., xsrl; falsification). For instance, xSRL xsrl combines local post-hoc explanations with global abstract policy graphs to identify and patch safety violations. AEGIS-RL falsification builds on xSRL by developing a verification approach that integrates formal model checking with risk- and uncertainty-guided falsification. However, repurposing these single-agent methods for MARL is not straightforward due to the heightened complexity of multi-agent systems. The explainability challenge in MARL surpasses that of single-agent RL, primarily due to the inherent coupling and dependencies among agents. While single-agent explanations can focus on an individual policy, MARL requires untangling a web of interdependencies, compounded by emergent collective behaviors and environmental non-stationarity. These dynamic feedback loops obscure cause-effect relationships, making it difficult to distinguish between agent-level and system-level failures.
@@
Moreover, conventional failure analysis in MARL often tracks global metrics such as team reward or derived statistics over time performance1; performance2. Such global performance metrics can obscure policy instability in individual agents Abdallah2009WhyGP; performance3 as we show in our experiments. Moreover, they neither identify the first agent to enter a non-robust state nor reveal how deviations propagate, all of which are needed to pinpoint weak points and pathways for proactive mitigation.
@@
Therefore, this paper formalizes the interpretable failure analysis problem in MARL around three core questions; Q1. Who is the true Patient-0: the first agent to enter a non-robust state?, Q2. Why might a non-attacked agent be flagged first? How do we distinguish true sources from amplified downstream effects?, and Q3. How does instability propagate across agents over time through the system’s learned coordination pathways? We propose to address these questions through a two-stage, gradient-based forensic method. Stage 1 performs per-agent local detection by probing the Taylor-error of a policy-gradient cost; the earliest threshold crossing declares an initial Patient-0 candidate (Q1). Stage 2 validates or corrects this candidate by tracing accelerating upstream influence using critic derivatives-first- and second-order action sensitivities aggregated over causal windows (Q2). This approach naturally yields constructions of directed contagion graphs that summarize propagation pathways, influence strength, and timing (Q3).
@@ 2. Related Work
Explainability in MARL (xMARL). Research in explainable MARL is limited and primarily splits into (i) credit assignment shap; shap2 and (ii) specification- or query-driven explanations temporal. Shapley-based methods shap; shap2 quantify each agent’s contribution to the team reward, offering insights into fairness and cooperation but not into failure dynamics or inter-agent cascades . Recent work continues to advance credit assignment, such as Multi-level Advantage Credit Assignment (MACA) MACA, which performs counterfactual reasoning across different levels of agent coordination but remains focused on optimizing training rather than diagnosing post-hoc failures . Complementarily, policy-level contrastive explanations temporal verify temporal task queries (e.g., encoded in PCTL*) via probabilistic model checking on an abstracted policy . While both lines illuminate aspects of MARL behavior, neither targets post-hoc failure analysis under compromised agents: they do not identify the entry point of a failure nor reconstruct how influence propagates across agents over time. Our method addresses this gap by detecting the first non-robust agent and mapping directed, time-resolved propagation paths using centralized-critic sensitivities.
Single-agent XRL and Policy Summaries. A rich body of XRL provides explanations for a single agent, including policy summarization via abstract policy graphs summarize; caps_aamas22, contrastive explaination methods contrastive, and frameworks that integrate local and global summaries xsrl; falsification. Safety-focused XRL (e.g., xSRL xsrl) ties explanations to debugging and patching unsafe behavior but remains single-agent in scope. Directly porting these methods to MARL misses the core challenge: the coupling and non-stationarity among agents obscure cause-effect relations across the team. Our work departs by explicitly modeling cross-agent influence during failure episodes and delivering a cascade map that explains how a local deviation spreads at run time.
Attacks and Robustness. The vulnerability of cooperative MARL to adversarial pressure is well-studied, including attacks that degrade team performance MARLPrima; ijcai23_anomaly; ami_attack; Lischke. These works establish vulnerabilities and propose defenses, but generally focus on performance degradation or anomalous behavior detection without explaining the failure’s origin and propagation pathway. Our contribution is orthogonal and complementary: a post-hoc explainer that identifies Patient-0 and attributes influence across agents.
@@ 3. Approach
We propose a two-stage, interpretable gradient-based method for MARL failure forensics. The key insight is that a failure typically begins with one agent and then propagates via learned coordination. Stage 1 performs per-agent local detection by probing policy-gradient Taylor error and declares Patient-0 as the first threshold crossing. Stage 2 validates or corrects it by tracing accelerating upstream influence using critic derivatives-first-order sensitivity and directional curvature-aggregated over a short causal window. These scores yield a directed contagion graph that summarizes who influenced whom, the strength and frequency of amplification, and how quickly deviations spread. We assume a single failure entry per episode, with subsequent failures arising from cascades through cooperative coupling. [^1] Relaxing the single–entry assumption is left to future work.
@@ 3.1. Two-Stage Patient-0 Detection and Validation
so that for small $\eta_{i}$, $\mathcal{L}_{i}^{t}(\eta_{i})\approx\tfrac{1}{2}\eta_{i}^{\top}\nabla_{o_{i}}^{2}J_{i}(o_{i}^{t},\tau_{i})\eta_{i}$. This provides a curvature probe in observation space without forming the full Hessian, similar to the Second-Order Identification of Non-Robust Directions test from single-agent RL soinrd. Large deviations indicate the policy is in a non-robust, high-curvature region where small observation changes could cause large policy shifts.
@@
In MARL algorithms with action-value critic (e.g., MADDPG lowe2017multi), we leverage the trained per-agent critic $Q_{i}(s,a_{1},\dots,a_{n})$ that conditions on the global state (or joint observations) and the joint action; this object encodes inter-agent dependencies learned during training. In algorithms that only learn a state-value function (e.g., HATRPO hatrpo), we fit a light post-hoc probe critic $\tilde{Q}_{i}(s,a_{1},\dots,a_{n})$ on frozen rollouts for attribution only (not for control), by regressing to TD or Monte-Carlo targets. [^2] Either critic is required only to be differentiable with respect to each action argument so that we can take first- and second-order derivatives. For discrete actions we use a standard continuous relaxation (e.g., softmax with temperature or Gumbel–Softmax during probing); for continuous actions we directly backpropagate. This yields well-defined partials $\partial Q_{i}/\partial a_{j}$ and $\partial^{2}Q_{i}/\partial a_{j}^{2}$ that quantify how agent $j$’s action locally perturbs agent $i$’s value and, by extension, its policy’s stability signals. We define the critic cost function as $L_{Q_{i}}=-Q_{i}$, thereby treating higher values as worse and lower values as better, which preserves the standard characteristics of a cost function.
@@ 4.1. Setup
Environments and algorithms. We use Simple Spread (cooperative navigation) with $n\!\in\!\{3,5\}$ agents and SMAC (Starcraft II). In Simple Spread, $n$ agents must navigate to $L$ landmarks through coordinated movement. Agents observe relative positions of other agents and landmarks, and receive shared rewards to minimize distance to landmarks. SMAC is a decentralized micromanagement benchmark where 3 Marine agents must defeat 3 Zealot opponents through coordinated focus fire and positioning. Each agent receives partial observations and controls one Marine unit, requiring emergent coordination strategies to overcome the enemy’s statistical advantages. We tested two MARL algorithms; MADDPG lowe2017multi and HATRPO hatrpo.
@@
Failure Simulation/ Intervention Protocol. We simulate failures using the worst-action attack from adversarial MARL ami_attack; MARLPrima. We then created an intervention protocol to test our proposed work. For each ordered pair $(j\!\to\!i)$, we intervene at two classes of timesteps computed from the trained critic(s): Critical states where $G_{ij}$ is high and $D_{ij}\!>\!0$, and Robust states where $G_{ij}$ is low and $D_{ij}\!\leq\!0$. We replace $a_{j}$ with the same-strength attack action in both conditions and compare outcomes over an identical window $\mathcal{W}$. Per environment this yields $500$ base episodes $\times$ ordered pairs $\times 2$ conditions: 6,000 variants for n=3n{=}3 (6 pairs), and 20,000 for n=5n{=}5 (20 pairs). For the SMAC 3s_vs_3z, we experimented with $100$ base episodes, yielding 1,200 variants in total.
@@
Evaluation Metrics. To validate (Patient-0 identification), we report four accuracy measures: Stage-1 accuracy $\Pr(\hat{p}_{\mathrm{S1}}{=}p^{\star})$, Traceback accuracy $\%(\text{chain ends at }p^{\star})$, Combined accuracy after Stage-2 correction, and Correction rate (the fraction of downstream-first misidentifications successfully corrected by Stage-2, calculated over all downstream-first cases). We didn’t report Patient-0 baselines because prior works address different problems: Lischke detects per-timestep anomalies using LSTMs on state-action traces without identifying the initial failure source, while ijcai2023p19 develops decentralized GRU-based normality scoring for anomaly detection but does not identify the first inciting agent.
@@ 4.2. Results
Case 2 - Downstream-first flag with traceback correction (Simple Spread, n=5n{=}5, MADDPG). An adversarial perturbation is applied to agent 3 over $t{=}5\text{--}8$. Yet Stage 1 raises the first alarm at agent 1 at t=7t{=}7, creating a downstream-first cascade and a false Patient-0.